# Who’s the GOAT? Sports Rankings and Data-Driven Random Walks on the Symmetric Group

### v1

@@ Abstract.
Given a collection of historical sports rankings, can one tell which player is the greatest of all time (i.e., the GOAT)? In this work, we design a data-driven random walk on the symmetric group to obtain a stationary distribution over player rankings, spanning across different time periods in sports history. We combine this distribution with a notion of stochastic dominance to obtain a partial order over the players. We implement our methods using publicly available data from the Association of Tennis Professionals (ATP) and the Women’s Tennis Association (WTA) to find the GOATs in the respective categories.
@@ 1. Introduction
For instance, consider Steffi Graf and Serena Williams; two widely successful players in women’s professional tennis. Between the two, Williams leads in the number of major titles, holding 23 compared to the 22 held by Graf. However, Graf leads in the number of titles, holding 107 compared to the 73 held by Williams. Which of these statistics—the number of titles or the number of major titles—is more important? How does $1$ additional major title fare against $34$ additional titles? Note that these players had little overlap in competition: Graf played professionally from 1982 to 1999, whereas Williams turned professional in 1995 and only retired recently in 2022. How do the numerous differences in professional tennis from the 1980s to the 2020s (e.g., changes in equipment, training, court surfaces, and point systems; consolidation of prestige around certain tournaments; increased competition/access to the sport; a global pandemic that suspended multiple tournaments in 2020) relate to the significance of these and other performance statistics?
@@ 1.1. Contributions
In this work, we propose a mathematical framework that leverages historical ranking data to compare players, and in particular players spanning across different time periods in sports history. In broad strokes, our method is as follows:
@@
    First, we design a data-driven random walk on the symmetric group to obtain a stationary distribution over player rankings.
@@
    Next, we combine this distribution with a notion of stochastic dominance to obtain a partially ordered set (poset) of players.
This approach to player comparison represents a departure from existing methods in two principal ways. First, compared to the use of individual performance statistics or match-level data, the use of historical rankings more directly reflects the evolution of circumstances and/or value systems. For example, players might be strategic around tournament participation in response to the ranking system in place—what one had to do to be top-ranked in the 1980s might be different from what one had to do in the 2010s—and this might not be as evident in other datasets. Moreover, it opens the opportunity for longitudinal comparisons in ranked systems for which match-level data is not available/does not exist (e.g., individual players within team sports, music charts, college rankings). Second, the adoption of a notion of stochastic dominance helps come to terms with the possibility that there is no single GOAT, instead limiting pairwise comparisons to cases in which these can be established conclusively.
To illustrate the practical merits of our approach, we perform a numerical implementation using singles ranking data from the “Open Era” of the ATP and the WTA. We experiment with a concept of GOAT that captures individual performance at the top of ranking cutoffs—performance as a “top 3,” “top 5,” “top 10,” or “top 20” player (refer to Section 3.1 for details)—to obtain data-driven posets and find the GOATs in the respective categories.
@@
- Federer and Nadal better than Sampras
- Djokovic and Sampras incomparable
@@
- All better than Connors, Lendl, and Sampras
- Federer and Nadal better than Becker
- Djokovic and Becker incomparable
Top 20 Federer and Nadal
- Federer better than Djokovic
- Nadal and Djokovic incomparable
@@
- Graf better than Navratilova
@@
- Graf and S. Williams better than Seles
@@
- Navratilova better than Seles
- Graf and Seles incomparable
@@
Graf, Navratilova, Seles, S. Williams,
@@ V. Williams, and Wozniacki
Table 2. Summary of GOATs of the WTA. The surnames mentioned correspond to Steffi Graf, Martina Navratilova, Monica Seles, Serena Williams, Venus Williams, and Caroline Wozniacki (in alphabetical order). Refer to Section 4.3 for more details.
@@ 1.2. Related Work
Radicchi [radicchi2011best] addressed the question of who is the GOAT in men’s professional tennis using historical match-level data and a network diffucion algorithm similar to PageRank [brin1998anatomy]. Baker and Hale [baker2014dynamic] addressed the same question using time-varying paired comparison models and the results of matches in major tournaments. Temesi, Szádoczki, and Bozóki [temesi2024incomplete] addressed the question of who is the GOAT in women’s professional tennis using pairwise comparison matrices and the results of matches between players who were ever ranked first. Still, from a philosophical point of view, sports rankings have been criticized for their inability to capture comparisons that are “too close to call” [bordner2016all].
In contrast to existing methods, we propose the aggregation of historical rankings into a single poset as an alternative to the stringent requirements of a linear order. That is, some pairs of players can be conclusively compared, others cannot, and there might be various, pairwise incomparable GOATs. Thus, our conceptualization of the problem is most closely related to that of rank aggregation [dwork2001rank, schalekamp2009rank] (refer also to Lin [lin2010rank] for a survey).
However, our methods are actually most closely related to those from the theory of card shuffling—the design and analysis of methods to sample uniform random permutations. The set $[n]\coloneqq\{1,2,\ldots,n\}$ corresponds to a deck of $n\in\mathbb{N}\coloneqq\{1,2,\ldots\}$ cards, and different card shuffling procedures (e.g., riffle shuffling, shuffling by transposing two random cards, shuffling by transposing two random adjacent cards) correspond to different sampling methods. The central questions in this area relate to convergence rates: how many times does a given shuffling procedure need to be repeated so that a given deck is sufficiently close to uniformly distributed? Bayer and Diaconis [bayer1992trailing] showed that $(3/2)\log n+\epsilon$ riffle shuffles are necessary and sufficient for a deck of $n$ cards to start becoming well-mixed, exhibiting a cutoff phenomenon [aldous1986shuffling] (so that a standard deck of $52$ cars requires about $6$ repetitions). Diaconis and Shahshahani [diaconis1981generating] showed that the random transposition shuffle has a cutoff at $(1/2)n\log n$, whereas Lacoin [lacoin2016mixing] showed that the random adjacent transposition shuffle has a cutoff at $(1/\pi^{2})n^{3}\log n$. Refer to Diaconis and Fulman [diaconis2023mathematics] for more on card shuffling.
@@ 1.3. Organization
The remainder of this work is organized as follows. For the purpose of self-containment, in Section 2 we summarize some necessary technical background. In Section 3 we design our methods. In Section 3 we implement our methods using data from the ATP and the WTA, and analyze the results of our numerical implementation. Lastly, in Section 5 we make concluding remarks.
@@ 2.1. Symmetric Group
Refer to any abstract algebra textbook for more on the symmetric group.
@@ 2.2. Markov Chains
for all $t\in\mathbb{N}$ whenever the conditional probabilities are well-defined. If the random variables share a finite co-domain, then the chain has a finite state space—say, indexed by $[m]$ for some $m\in\mathbb{N}$. The chain is time-homogeneous if
@@ 2.3. Sports Rankings and Partially Ordered Sets
A partially ordered set (poset) is a pair $\pi=([n],\preceq_{\pi})$ consisting of a ground set $[n]$ and a partial order $\preceq_{\pi}$ satisfying:
@@
A poset may be represented pictorially by a Hasse diagram, which is a drawing of its transitive reduction. For example, if the edges in the graphs in Figure 1 were all oriented upwards, we would obtain Hasse diagrams of two different posets on the symmetric group of order $4$, both with unique minimal element $1234$ and unique maximal element $4321$.
@@
Let $[n]$ represent a set of players. A ranking is a totally ordered set $\pi=([n],\preceq_{\pi})$, which we treat as a permutation $\pi\in\mathfrak{S}_{n}$ (in explicit notation) given the natural ordering $\leq$ of $\mathbb{N}$. In particular, we have that
@@
of historical sports rankings on subsets of players. In subsequent sections we will consider partial orders on the full set of players $[n]$ together with their corresponding linear extensions as part of our problem output.
@@ 3.1. Cutoffs and the meaning of GOAT
In the experimental part of this work, we restrict the rankings in $\mathcal{R}$ to involve the top $3$, $5$, $10$, and $20$ players as cutoffs from the full ranking data that might be available. We do this for two reasons:
@@ 3.2. Transition Rule
for all distinct $i,j\in[n]$. The second term in the product, |$\left|\{r\in\mathcal{R}_{ij}:j\prec_{r}i\}\right|$, encodes the number of (cut off) historical rankings in which player $i$ is ranked better than player $j$. We scale this quantity by the first term in the product, $\left|\mathcal{R}_{i}\right|/\left|\mathcal{R}_{j}\right|$, which is the players’ relative prevalence above the cutoff. The motivation for this is as follows.
@@
Suppose players $i$ and $j$ have short and long prevalence in the (cut off) historical rankings, respectively. In particular, it might be that player $i$ made it to the top of the ranking for a brief period of time, consequently beating player $j$ during this period. However, since the prevalence of player $i$ is short, there might be significant periods of time during which player $i$ fell below the cutoff while player $j$ did not—by definition, player $j$ beat player $i$ during these periods. This artifact of cutoffs is not captured by the term |$\left|\{r\in\mathcal{R}_{ij}:j\prec_{r}i\}\right|$, which by itself favors player $i$ over player $j$. We correct for this in the model using their prevalence ratio $\left|\mathcal{R}_{i}\right|/\left|\mathcal{R}_{j}\right|$.
@@ 3.3. Partial Ordering
Let $\mathcal{L}(\preceq_{\mu})$ denote the set of linear extensions of $\preceq_{\mu}$ (refer to Section 2.3). In the experimental part of this work, we will consider the average rank of a player with respect to a random linear extension as a summary statistic of the player’s career performance. In particular, for each $i\in[n]$ we will consider
$\bar{X}(i)=\frac{1}{N}\sum_{t=1}^{T}X_{t}(i),$ (5)
@@ 3.4. Sampling
Given the sports analytics scope of this work, we do not attempt to answer these questions and simply acknowledge that analytical questions around data-driven random walks on the symmetric group are of independent mathematical interest. In the experimental part of this work, we execute this random walk for a large number of steps collecting every $2(n-1)$th sample. Given the $n-1$ diameter noted in Remark 2.6 and the $1/2$ probability of taking a lazy step in (3), this is the expected number of steps needed for the entire state space to be reachable from any given state.
@@ 4. Numerical Implementation
In this section, we implement our methods using singles rankings from the “Open Era” of the ATP and the WTA. As discussed in Section 3.1, we experiment with taking the top $3$, $5$, $10$, and $20$ players as different cutoffs from the full ranking data that is available. For each experiment, we execute a random walk as described in Section 3.4 until we collect $100,000$ samples. We then use these samples to obtain data-driven posets over the players, as described in Section 3.3. We illustrate these posets as horizontal Hasse diagrams with maximal elements on the right-hand side (the full diagrams can be found in the Appendix). To estimate the average rank of the players in a given poset, we first sample $100,000$ linear extensions uniformly at random using SageMath [sagemath] and then follow (5).
The remainder of this section is organized as follows. In Section 4.1 we further describe the datasets we use for our experiments. In Section 4.2 we present our results using rankings from the ATP, whereas in Section 4.3 we present our results using rankings from the WTA. Refer to Tables 1 and 2 for a summary of some of the results. Lastly, in Section 4.4 we compare the results obtained for the ATP and the WTA to note qualitative differences between the two associations. All mentioned statistics that are external to our numerical implementation can be verified online [ATP_Tour_2024, WTA_Tennis_2024].
@@ 4.2. Rankings from the ATP
In Figure 2 we plot the top portion of the Hasse diagram of the poset over “top 3 players,” where Djokovic, Federer, Nadal, and Sampras are incomparable maximal elements—no one of these players is conclusively better than another. Note also that they all have a similar average rank. Conversely, we conclude that each of Djokovic, Federer, Nadal, and Sampras is better than Connors and Lendl as a top 3 player. Similarly, as top 3 players, Connors is better than Andre Agassi, who is better than Björn Borg, who is better than Andy Murray, who is better than Carlos Alcaraz. In other words, these players form a chain in the poset. Formally, this is
@@
In Figure 3 we plot the top part of the Hasse diagram of the poset over “top 5 players,” where Djokovic, Federer and Nadal remain incomparable maximal elements, although Federer fares slightly better than the other two in terms of average rank. However, Sampras is no longer a maximal element: as a top 5 player, he is worse than both Federer and Nadal, and he is incomparable to Djokovic. Sampras is moreover incomparable to Connors and Lendl (both of whom he was better than as a top 3 player). Note that Djokovic and Connors are incomparable as top 5 players: their difference in average rank favors Djokovic, but it is not sufficiently conclusive. Moreover, note that Alcaraz no longer appears near the top of the Hasse diagram, having been displaced by other players who are better top 5 players than they are top 3 players, such as Guillermo Vilas (this is reflected in the average rank plot). First, note that Alcaraz and Vilas did not overlap in competition: Vilas played professionally from 1968 to 1992. Second, note that Alcaraz transitioned very quickly from below the top 20 to within the top 3 from February to September of 2022, and has consistently prevailed within the top 3 until the time of writing. His prevalence within the top 3 is significant independently of his short career at present. However, when the cutoff is relaxed to top 5, his short career-to-date puts him at a disadvantage compared to retired players with significant career-long prevalence in the top 5 (but perhaps not in the top 3), such as Vilas.
In Figure 4 we plot the top part of the Hasse diagram of the poset over “top 10 players,” where Djokovic, Federer, and Nadal once again remain incomparable maximal elements. However, unlike the previous cutoff, we conclude that Djokovic is better than Connors as a top 10 player. Qualitatively, we observe the inclusion of more players near the top of this Hasse diagram compared to the previous cases, indicating that there are more “good top 10 players” than there are good top 3 or top 5 players. For example, Becker (who was far from maximal in Figures 2 and  3) is now close to maximal: only Federer and Nadal are better than him, and he is incomparable to Djokovic.
@@
Lastly, in Figure 5 we plot the top part of the Hasse diagram of the poset over “top 20 players.” Interestingly, only Federer and Nadal remain incomparable maximal elements. Despite this, only Federer is conclusively better than Djokovic as a top 20 player—Djokovic and Nadal are incomparable as top 20 players. Qualitatively, we again observe the inclusion of even more players near the top of this Hasse diagram compared to the previous cases. For instance, the Djokovic $\succ_{\mu}$ Connors $\succ_{\mu}$ Agassi chain from Figure 2 is broken in Figure 5, where all three of these players are equally close to the top of the Hasse diagram and form an anti-chain.

### v2

@@ Abstract.
Given a collection of historical sports rankings, can one tell which player is the greatest of all time (i.e., the GOAT)? In this work, we design a data-driven random walk on the symmetric group to obtain a stationary distribution over player rankings, spanning across different time periods in sports history. We combine this distribution with a notion of stochastic dominance to obtain a partial order over the players. Compared to existing methods, our approach is distinct in that i) using historical rankings captures the evolution of value systems and facilitates player comparisons when head-to-head data is unavailable, and ii) aggregating into a partial order formally comes to terms with the possibility that, while some player comparisons can be established conclusively, others are “too close to call.” We implement our methods using publicly available data from the Association of Tennis Professionals (ATP) and the Women’s Tennis Association (WTA). Our main findings indicate that Steffi Graf and Serena Williams are the ones that come ahead as the GOATs of the WTA. Likewise, the “Big Three,” that is Novak Djokovic, Roger Federer, and Rafael Nadal, are the ones that come ahead as the GOATs of the ATP. As a secondary finding, we note major differences in terms of career and dominance longevity for top players across the associations. While initially motivated by this application in sports analytics, our methods can also be applied to other practical domains where deriving rankings from historical data can inform operational decisions, such as route planning logistics.
@@ 1. Introduction
For instance, consider Steffi Graf and Serena Williams; two widely successful players in women’s professional tennis. Between the two, Williams leads in the number of major titles, holding 23 compared to the 22 held by Graf. However, Graf leads in the overall number of titles, holding 107 compared to the 73 held by Williams. Which of these statistics—the number of titles or the number of major titles—is more important? How does $1$ additional major title fare against $34$ additional titles? Note that these players had little overlap in competition: Graf played professionally from 1982 to 1999, whereas Williams turned professional in 1995 and only retired recently in 2022. How do the numerous differences in professional tennis from the 1980s to the 2020s (e.g., changes in equipment, training, court surfaces, and point systems; consolidation of prestige around certain tournaments; increased competition/access to the sport; a global pandemic that suspended multiple tournaments in 2020) relate to the significance of these and other performance statistics?
@@ 1.1. Contributions
In this work, we propose a mathematical framework that leverages historical ranking data to compare players, and in particular players spanning across different time periods in sports history. We summarize our method in Figure 1. In broad strokes, it is as follows:
@@
    First, we use cutoffs of historical ranking data (Figure 1(a)) to design a data-driven random walk on the symmetric group, from which we obtain a stationary distribution over rankings on the full set of players (Figure 1(b)).
@@
    Next, we combine this distribution with a notion of stochastic dominance to obtain a partially ordered set (poset) of players (Figure 1(c)).
(a) Historical rankings.
(b) Data-driven random walk.
(c) Poset over players.
Figure 1. Outline of our method. Figure 1(a): The input data consists of the top portions of the full historical rankings, at varying cutoffs. Figure 1(b): The random walk is designed to have a stationary distribution over rankings on the full set of players, reflecting patterns of relative order present in the input data. Figure 1(c): The output data consists of a poset over the full set of players, based on a notion of stochastic dominance.
This approach to player comparison represents a departure from existing methods in two principal ways. First, compared to the use of individual performance statistics or match-level data, the use of historical rankings more directly reflects the evolution of circumstances and/or value systems. For example, players might be strategic around tournament participation in response to the ranking system in place—what one had to do to be top-ranked in the 1980s might be different from what one had to do in the 2010s—and this might not be as evident in other datasets. Moreover, it opens the opportunity for longitudinal comparisons in ranked systems for which match-level data is not available/does not exist (e.g., individual players within team sports, music charts, college rankings). Second, the adoption of a notion of stochastic dominance helps formally come to terms with the possibility that there is no single GOAT, instead limiting pairwise comparisons to cases in which these can be established conclusively.
To illustrate the practical merits of our approach, we perform a numerical implementation using singles ranking data from the “Open Era” of the ATP and the WTA. We experiment with a concept of GOAT that captures individual performance at the top of ranking cutoffs—performance as a “top 3,” “top 5,” “top 10,” or “top 20” player (refer to Figure 1(a) and Section 3.1 for details)—to obtain data-driven posets and find the GOATs in the respective categories.
@@
- Federer and Nadal better than Sampras and Connors
- Connors, Djokovic, Lendl, and Sampras incomparable
@@
- All better than Connors and Sampras
- Federer and Nadal better than Becker and Lendl
- Becker, Djokovic, and Lendl incomparable
Top 20 Federer, Nadal, and Lendl
- All better than Djokovic
- Djokovic better than Sampras
@@
- Graf better than Hingis and Navratilova
- S. Williams better than Hingis
@@
- Graf and S. Williams better than Seles and Hingis
@@
- Graf and Navratilova better than Seles
@@
Graf, Navratilova, S. Williams,
@@ V. Williams, and Wozniacki
Table 2. Summary of GOATs of the WTA. The surnames mentioned correspond to Steffi Graf, Martina Hingis, Martina Navratilova, Monica Seles, Serena Williams, Venus Williams, and Caroline Wozniacki (in alphabetical order). Refer to Section 4.3 for more details.
@@ 1.2. Related Work
Radicchi [radicchi2011best] addressed the question of who is the GOAT in men’s professional tennis using historical match-level data and a network diffucion algorithm similar to PageRank [brin1998anatomy]. Baker and McHale [baker2014dynamic] addressed the same question using time-varying paired comparison models and the results of matches in major tournaments. Temesi, Szádoczki, and Bozóki [temesi2024incomplete] addressed the question of who is the GOAT in women’s professional tennis using pairwise comparison matrices and the results of matches between players who were ever ranked first. Still, from a philosophical point of view, sports rankings have been criticized for their inability to capture comparisons that are “too close to call” [bordner2016all].
In contrast to existing methods, we propose the aggregation of historical rankings into a poset as an alternative to the stringent requirements of a linear order. That is, some pairs of players can be conclusively compared, others cannot, and there might be various, pairwise incomparable GOATs. Thus, our conceptualization of the problem is most closely related to that of rank aggregation [dwork2001rank, schalekamp2009rank] (refer also to Lin [lin2010rank] for a survey).
However, our methods are actually most closely related to those from the theory of card shuffling—the design and analysis of methods to sample uniform random permutations of a fixed finite set. The set $[n]\coloneqq\{1,2,\ldots,n\}$ corresponds to a deck of $n\in\mathbb{N}\coloneqq\{1,2,\ldots\}$ cards, and different card shuffling procedures (e.g., riffle shuffling, shuffling by transposing two random cards, shuffling by transposing two random adjacent cards) correspond to different sampling methods. The central questions in this area relate to convergence rates: how many times does a given shuffling procedure need to be repeated so that a given deck is sufficiently close to uniformly distributed? Bayer and Diaconis [bayer1992trailing] showed that $(3/2)\log n+\epsilon$ riffle shuffles are necessary and sufficient for a deck of $n$ cards to start becoming well-mixed, exhibiting a cutoff phenomenon [aldous1986shuffling] (so that a standard deck of $52$ cars requires about $6$ repetitions). Diaconis and Shahshahani [diaconis1981generating] showed that the random transposition shuffle has a cutoff at $(1/2)n\log n$, whereas Lacoin [lacoin2016mixing] showed that the random adjacent transposition shuffle has a cutoff at $(1/\pi^{2})n^{3}\log n$. Refer to Diaconis and Fulman [diaconis2023mathematics] for more on card shuffling.
@@ 1.3. Organization
The remainder of this work is organized as follows. For the purpose of self-containment, in Section 2 we summarize some necessary technical background. In Section 3 we design our methods. In Section 3 we implement our methods using data from the ATP and the WTA, and analyze the results of our numerical implementation. Lastly, in Section 5 we make concluding remarks and point out other potential applications of our methods.
@@ 2.1. Symmetric Group
Refer to any introductory abstract algebra textbook for more on the symmetric group.
@@ 2.2. Markov Chains
for all $t\in\mathbb{N}$ whenever the conditional probabilities are well-defined. If the random variables share a finite co-domain, then the chain has a finite state space. Say, it is indexed by $[m]$ for some $m\in\mathbb{N}$. The chain is time-homogeneous if
@@ 2.3. Sports Rankings and Partially Ordered Sets
A finite partially ordered set (poset) is a pair $\pi=([n],\preceq_{\pi})$ consisting of a ground set $[n]$ and a relation $\preceq_{\pi}$ satisfying:
@@
A poset may be represented pictorially by a Hasse diagram, which is a drawing of its transitive reduction. Formally, for any $i,j\in[n]$, we say that $j$ covers $i$ if $i\prec_{\pi}j$ and there is no $k\in[n]$ satisfying $i\prec_{\pi}k\prec_{\pi}j$. Then, the Hasse diagram of $\pi$ is a directed graph whose vertices are $[n]$ and whose edges are given by the cover relations. For example, if the edges in the graphs in Figure 2 were all oriented upwards, we would obtain Hasse diagrams of two different posets on the symmetric group of order $4$, both with unique minimal element $1234$ and unique maximal element $4321$.
@@
Now, let $[n]$ represent a set of players. A ranking is a totally ordered set $\pi=([n],\preceq_{\pi})$, which we treat as a permutation $\pi\in\mathfrak{S}_{n}$ (in explicit notation) given the natural ordering $\leq$ of $\mathbb{N}$. In particular, we have that
@@
of historical sports rankings on subsets of players (refer to Figure 1(a)). In subsequent sections we will consider partial orders on the full set of players $[n]$ together with their corresponding linear extensions as part of our problem output (refer to Figure 1(c)).
@@ 3.1. Cutoffs and the meaning of GOAT
In the experimental part of this work, we restrict the rankings in $\mathcal{R}$ to involve the top $3$, $5$, $10$, and $20$ players as cutoffs from the full ranking data that might be available. This is illustrated in Figure 1(a). We do this for two reasons:
@@ 3.2. Transition Rule
for all distinct $i,j\in[n]$. The second term in the product, |$\left|\{r\in\mathcal{R}_{ij}:j\prec_{r}i\}\right|$, encodes the number of cut off historical rankings in which player $i$ is ranked better than player $j$. We scale this quantity by the first term in the product, $\left|\mathcal{R}_{i}\right|/\left|\mathcal{R}_{j}\right|$, which is the players’ relative prevalence above the cutoff. The motivation for this is as follows.
@@
Suppose players $i$ and $j$ have short and long prevalence in the cut off historical rankings, respectively. In particular, it might be that player $i$ made it to the top of the ranking for a brief period of time, consequently beating player $j$ during this period. However, since the prevalence of player $i$ is short, there might be significant periods of time during which player $i$ fell below the cutoff while player $j$ did not—by definition, player $j$ beat player $i$ during these periods. This artifact of cutoffs is not captured by the term |$\left|\{r\in\mathcal{R}_{ij}:j\prec_{r}i\}\right|$, which by itself favors player $i$ over player $j$. We correct for this in the model using their prevalence ratio $\left|\mathcal{R}_{i}\right|/\left|\mathcal{R}_{j}\right|$.
@@ 3.3. Partial Ordering
Let $\mathcal{L}(\preceq_{\mu})$ denote the set of linear extensions of $\preceq_{\mu}$. In the experimental part of this work, we will consider the average rank of a player with respect to a random linear extension of $\preceq_{\mu}$ as a summary statistic of the player’s historical performance. In particular, for each $i\in[n]$ we will consider
$\overline{X}(i)=\frac{1}{N}\sum_{t=1}^{T}X_{t}(i),$ (5)
@@ 3.4. Sampling
Given the applied scope of this work, we do not attempt to answer these questions and simply acknowledge that analytical questions around data-driven random walks on the symmetric group are both well-motivated and of independent mathematical interest. In the experimental part of this work, we execute this random walk for a large number of steps collecting every $2(n-1)$th sample. Given the $n-1$ diameter noted in Remark 2.6 and the $1/2$ probability of taking a lazy step in (3), this is the expected number of steps needed for the entire state space to be reachable from any given state.
@@ 4. Numerical Implementation
In this section, we implement our methods using singles rankings from the “Open Era” of the ATP and the WTA. As discussed in Section 3.1, we experiment with taking the top $3$, $5$, $10$, and $20$ players as different cutoffs from the full ranking data that is available. For each experiment, we execute a random walk as described in Section 3.4 until we collect $100,000$ samples. We then use these samples to obtain data-driven posets over the players, as described in Section 3.3. We illustrate these posets as horizontal Hasse diagrams with maximal elements on the right-hand side. The full Hasse diagrams we obtain are shown in Figure 3, but in short we will zoom in on their top portions.
(a) Rankings from the ATP.
(b) Rankings from the WTA.
To estimate the average rank of the players in a given poset, we first sample $100,000$ linear extensions uniformly at random using the package SageMath by [sagemath] and then follow (5).
The remainder of this section is organized as follows. In Section 4.1 we further describe the datasets we use for our experiments. In Section 4.2 we discuss our results using rankings from the ATP, whereas in Section 4.3 we discuss our results using rankings from the WTA. Refer to Tables 1 and 2 for a summary of some of our findings. Lastly, in Section 4.4 we compare the results obtained for the ATP and the WTA to note qualitative differences between the two associations. All mentioned statistics that are external to our numerical implementation can be verified online [ATP_Tour_2024, WTA_Tennis_2024]. The source code for these experiments is available at github.com/jcmartinezmori/goat.
@@ 4.2. Rankings from the ATP
In Figure 4 we plot the top portion of the Hasse diagram of the poset over “top 3 players,” where Djokovic, Federer, Nadal, and Sampras are incomparable maximal elements—no one of these players is conclusively better than another. Note also that they all have a similar average rank. Conversely, we conclude that each of Djokovic, Federer, Nadal, and Sampras is better than Connors and Lendl as a top 3 player. Similarly, as a top 3 player, Connors is better than Andre Agassi, who is better than Björn Borg, who is better than Andy Murray, who is better than Carlos Alcaraz. In other words, these players form a chain in the poset. Formally, this is
@@
In Figure 5 we plot the top part of the Hasse diagram of the poset over “top 5 players,” where Djokovic, Federer and Nadal remain incomparable maximal elements, although Federer fares slightly better than the other two in terms of average rank. However, Sampras is no longer a maximal element: as a top 5 player, he is worse than both Federer and Nadal, and he is incomparable to Djokovic. Sampras is moreover incomparable to Connors and Lendl (both of whom he was better than as a top 3 player). Note that Connors, Djokovic, and Lendl are also incomparable as top 5 players: their difference in average rank favors Djokovic, but it is not sufficiently conclusive. Moreover, note that Alcaraz no longer appears near the top of the Hasse diagram, having been displaced by other players who are better top 5 players than they are top 3 players, such as Guillermo Vilas (this is reflected in the average rank plot). First, note that Alcaraz and Vilas did not overlap in competition: Vilas played professionally from 1968 to 1992. Second, note that Alcaraz transitioned very quickly from below the top 20 to within the top 3 from February to September of 2022, and has consistently prevailed within the top 3 until the time of writing. His prevalence within the top 3 is significant independently of his short career at present. However, when the cutoff is relaxed to top 5, his short career-to-date puts him at a disadvantage compared to retired players with significant career-long prevalence in the top 5 (but perhaps not in the top 3), such as Vilas.
@@
In Figure 6 we plot the top part of the Hasse diagram of the poset over “top 10 players,” where Djokovic, Federer, and Nadal once again remain incomparable maximal elements. However, unlike the previous cutoff, we conclude that Djokovic is better than Connors and Sampras as a top 10 player. Qualitatively, we observe the inclusion of more players near the top of this Hasse diagram compared to the previous cases, indicating that there are more “good top 10 players” than there are good top 3 or top 5 players. For example, Becker (who was far from maximal in Figures 4 and  5) is now close to maximal: only Federer and Nadal are better than him, and he is incomparable to Djokovic.
@@
Lastly, in Figure 7 we plot the top part of the Hasse diagram of the poset over “top 20 players.” Interestingly, Federer and Nadal are joined by Lendl as incomparable maximal elements. All three of them are better than Djokovic as a top 20 player. Qualitatively, we again observe the inclusion of even more players near the top of this Hasse diagram compared to the previous cases. For instance, the Djokovic $\succ_{\mu}$ Connors $\succ_{\mu}$ Agassi chain from Figure 4 is broken in Figure 7, where all three of these players are equally close to the top of the Hasse diagram and form an anti-chain.